OpenStudy (forach):

What is the frequency of a standing wave in a 1.5m closed pipe that is vibrating in the 5th harmonic if the speed of sound is 343m/s? A 285Hz B 390Hz C 572Hz D 620Hz

OpenStudy (niyex):

Simple Harmonic Motion \[f=\frac{ \lambda }{ v }\] there you go just input your value then you are good to go

OpenStudy (ijlal):

@niyex The formula is wrong actually the correct formula is \[v=f \lambda \] then, \[f=\frac{ v }{ \lambda }\] but for the 5th harmonic motion we now that \[\lambda=\frac{ 5v }{ 4L }\] Read up on the harmonic motion of waves in a pipe closed at one end. so we get our f to be \[f=\frac{ 5v }{ 4L }\] so we plug in the values we get \[f=\frac{ 5(343) }{ 4(1.5) }\] \[f=285.833 Hz\] that would make your answer to be Option A

OpenStudy (niyex):

thanks @ijlal. but i found something strange in one of your equation \[\lambda=\frac{ 5v }{ 4L } =f \frac{ 5v }{ 4L }\] are you saying that \[\lambda=f\] if that what your insinuating then bro please check your equations well cos frequency is quite different from wavelength. But i want to assume it is a mix up or a mistake

OpenStudy (ijlal):

@niyex oh sorry it was a typo :D thanks for pointing it out ! the equation is correct the mistake is \[\lambda = \frac{ 4L }{ 5 }\] so we put that in the equation \[f=\frac{ v }{ \lambda }\] it becomes \[f=\frac{ 5v }{ 4L }\]